show that 2 is a factor of n^2+3n
Answer by jim_thompson5910(35256)

n^2+3n = n(n+3)

Case 1: if n is even, then 2 is a factor of n(n+3) or n^2+3n (since n is a factor of n^2+3n)

Case 2: If n is odd, then n+3 is even (odd+odd = even), so 2 is factor of n+3 which means 2 is a factor of n^2+3n